OpenStudy (anonymous):

If a population of 1,000 individuals has 160 rr genotypes, what is the frequency of the rr genotype? Please explain how you got the answer.

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

1000 ----- 160 100 ----- x ----------------- x=(160*100)/1000 =16000/1000 =16 so this result mean that is 16%

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so just you need using the rule of 3 simples

OpenStudy (anonymous):

thanks so much! would you mind helping me with one other type of problem? @jhonyy9

OpenStudy (anonymous):

If a population of 1,000 individuals has 160 rr genotypes, assuming simple dominance by the A allele, what is the frequency of the dominant allele (R) in this population? Please explain? @jhonyy9

OpenStudy (anonymous):

@aaronq do you know this one?

OpenStudy (aaronq):

i dont remember how to do this sorry

OpenStudy (anonymous):

That's ok @aaronq

OpenStudy (anonymous):

@hba @jazy do either of you guys know how to do this? I'm so lost. My textbook makes it so confusing... I wish @jhonyy9 could explain this one again! :(

jhonyy9 (jhonyy9):

sorry but this one i dont undersatand it the first one of your question was easy but this last with A allele ??? sorry what mean this allele ? sorry but i dont understand it

OpenStudy (anonymous):

What do you mean? @jhonyy9

OpenStudy (anonymous):

A very large population of randomly mating laboratory mice contains 35% white mice. White coloring is caused by the double recessive genotype "aa." Calculate the frequency of the heterozygous genotype (Aa) in this population. Do you understand this one?

jhonyy9 (jhonyy9):

so can you rewrite me how makes your textbook ? why is it for you ,,confusing" ?

OpenStudy (anonymous):

I'm going to ask my teacher about it but she hasn't been answering... Do you understand how to do the other one I posted just now? @jhonyy9

jhonyy9 (jhonyy9):

moment please

jhonyy9 (jhonyy9):

100 ----- 35 white mice so what mean 35% ,,Aa" so do you know how many is this large population ?

OpenStudy (anonymous):

That was the only information given to me. When you write 100 --------- 35 does that mean you're multiplying?

jhonyy9 (jhonyy9):

no this you have wrote it above in your exercise 35% so this mean from 100 are white 35 ok ?

OpenStudy (anonymous):

I don't really understand the way you're typing... Can you try to word it differently? @jhonyy9

OpenStudy (anonymous):

My answer key says that the answer could be 47%, 48%, or 49% How did they get that?

jhonyy9 (jhonyy9):

so above you have wrote that 35%white mice is from a laboratory very large population and this is have wrote it that this 35% mean that from 100 mean 35 mice white ok ?

OpenStudy (anonymous):

Ok, I think I get that part. Now how do you get the answer of 47, 48, and 49 % ?

jhonyy9 (jhonyy9):

sorry but i not can getting these answers i dont know sure this way sorry

jhonyy9 (jhonyy9):

so this 49% if this is right --- so this mean approximatly a half part of this population

OpenStudy (anonymous):

yes, i think..

jhonyy9 (jhonyy9):

so than i dont know what is the way to get these answers you need to ask your teacher sorry

OpenStudy (anonymous):

Ok, can you try to help me with just one more? In corn, the allele for yellow coloring has complete dominance over the allele for white coloring. A random sample of 1,000 kernels from a population that is in equilibrium reveals that 910 are yellow and 90 are white. What is the percentage of heterozygotes in this population? The answer is .42 or 42. How did they get this??

jhonyy9 (jhonyy9):

sorry good luck bye

OpenStudy (anonymous):

Ok...